OpenStudy (anonymous):

How to graph this equation: -6/7x-1/2y=3/7. I'll give some of the options that I made up for you guys [As far as the graphs go]

OpenStudy (anonymous):

You need to add some more parentheses in order to make it more clear what this function is.

OpenStudy (anonymous):

There aren't any parentheses..

OpenStudy (anonymous):

Here's 1 option

OpenStudy (anonymous):

\[\frac{-6}{7x}-\frac{1}{2y}=\frac{3}{7}\] Is this the function?

OpenStudy (anonymous):

Like this.

OpenStudy (anonymous):

Well the function is confusing. How I would go to solving this problem is by isolating y in the function. Meaning that leave y by itself on one side and have everything on the other side. Then I would start plugging in random values of X such as x=0 and x=1 and see what I get for Y. Once I have enough points (X,Y) I can draw my own graph and see which one looks similar to the one I drew.

OpenStudy (anonymous):

Okay, thanks hun!(:

OpenStudy (anonymous):

I know it looks like a lot but this is the only way to solve it. Unless you can tell how the graph will be by just looking at the function which I can't even do.